1)Find the present values of these ordinary annuities. Discounting occurs once a year. Do not round intermediate calculations. Round your answers to the nearest cent. $200 per year for 16 years at 12%. $ $100 per year for 8 years at 6%. $ $1,000 per year for 8 years at 0%. $ Rework previous parts assuming they are annuities due. Present value of $200 per year for 16 years at 12%: $ Present value of $100 per year for 8 years at 6%: $ Present value of $1,000 per year for 8 years at 0%: $

2) What is the present value of a $200 perpetuity if the interest rate is 9%? If interest rates doubled to 18%, what would its present value be? Round your answers to the nearest cent. Present value at 9%: $ Present value at 18%: $

3) You borrow $110,000; the annual loan payments are $7,155.66 for 30 years. What interest rate are you being charged? Round your answer to the nearest whole number. %

4)Find the present values of the following cash flow streams at a 5% discount rate. Do not round intermediate calculations. Round your answers to the nearest cent. 0 1 2 3 4 5 Stream A $0 $150 $350 $350 $350 $300 Stream B $0 $300 $350 $350 $350 $150 Stream A: $ Stream B: $ What are the PVs of the streams at a 0% discount rate? Round your answers to the nearest dollar. Stream A: $ Stream B: $
Answer #1

1)

D 94.80 =PV(12%,16,-200,0,0) 620.98 =PV(6%,8,-100,0,0) $ 0.00 =PV(0%,8,-1000,0,0) 1 a) Present value of ordinary annuity 2 b)

*Please rate thumbs up